OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). a = 3b - 4 a + b =16

OpenStudy (anonymous):

(11,5)

OpenStudy (anonymous):

3b-4+b=16 4b-4 = 16 4b = 20 b = 5, a = 3*5-4 = 11

OpenStudy (anonymous):

Thanks